OpenStudy (anonymous):

The area of a given hexagon is equal to the area of an equilateral triangle whose perimeter is 36 inches. Find the length of a side of the regular hexagon.

OpenStudy (anonymous):

Each side of the triangle is 12 (36/3) Half of the "bottom", one of the sides, and the height of the triangle form another (right) triangle. The height of the triangle is sqrt(12^2-6^2) = sqrt(108) The area of the triangle is 6*sqrt(108)/2 The area of the hexagon is the same. A hexagon can be split into 6 equilateral triangles. Each triangle has an area of sqrt(108)/2 Now, reversing what I did earlier, we get sqrt(108)/2=y*sqrt((2y)^2-y^2)/2 Solve for y. sqrt(108)=y*sqrt((2y)^2-y^2) sqrt(108)=y*sqrt(4y^2-y^2) sqrt(108)=y*sqrt(y^2(4-1)) sqrt(108)=y*y*sqrt(3) sqrt(108)/sqrt(3)/y^2 y=sqrt(sqrt(108)/sqrt(3))